# ATK: Automatic Task-driven Keypoint Selection for Robust Policy Learning

### v1

@@ A.1 Mask network
To ensure the selection of the minimal set of $N$ keypoints, we enforce a sparsity penalty on the mask network. We denote the masking distribution over binary masks $m\in\{0,1\}^{M}$ for candidate keypoints $K$ as $q_{\phi}(m\mid K),$ and the action distribution conditioned on the masked keypoints as $\pi_{\theta}\bigl{(}a\mid m\odot K\bigr{)}.$ Because sampling $m\sim q_{\phi}(m\mid K)$ is discrete and non-differentiable, we employ the Gumbel–softmax relaxation [22] to enable gradient-based optimization. Concretely, a $K$-dimensional learnable parameters are used as logits for M independent Bernoulli probabilities, with each representing the likelihood of retaining the corresponding keypoint from the input candidate set. We then add Gumbel noise to each logit, divide by a temperature $\tau$, and apply softmax to form a continuous “soft" mask $y^{Soft}_{M}$. During the forward pass, we take $\arg\max_{i}\,y^{\mathrm{Soft}}_{i},$ to get a hard one-hot selection, but in the backward pass we use the smooth “soft" mask for gradient calculation. The policy $\pi_{\theta}(a\mid m\odot K)$ can be instantiated by any policy class model—e.g., a multi-layer perceptron, a Gaussian policy [45] or a score-based diffusion model [46]. The final objective jointly optimizes the policy $\pi_{\theta}(a\mid m\odot K)$ and the masking distribution $q_{\phi}(m\mid K)$, yielding a minimal, task-relevant keypoint representation.
@@ Appendix D Baseline implementation details
ResNet18 Encoder: We use a modified ResNet18 architecture as the encoder for depth inputs. The depth data is first resized to 224$\times$224 and duplicated across three channels to match the expected input format of the ResNet18 backbone. Then, we take the ResNet’s 512-dimensional feature embedding as a feature map and concatenate it with a 7-dimensional robot state (six joint values and a binary gripper state). The resulting vector is then passed to the diffusion policy [46] for action prediction.
R3M Encoder: To handle RGB inputs, we first resize the RGB images from 640$\times$480 resolution to 224$\times$224. Then, we use R3M[50] to extract a 512-dimensional feature embedding given image inputs and concatenate this embedding with the 7-dimensional joint and gripper state vector to form the input to the diffusion policy [46].
@@ E.1 Sim-to-real performance
Table 2: Real-world Policy Success Rates. Varying conditions including RP (random pose), RB (background), RO (distractor object), Light. $ATK$ consistently outperforms baseline methods using alternative modalities in sim-to-real transfer.
@@
FullSet 0.474$\pm$0.317 0.126$\pm$0.090 0.026$\pm$0.030 0.020$\pm$0.016 0.253$\pm$0.183 0.083$\pm$0.880 0.010$\pm$0.014 0.010$\pm$0.014
RandomSelect 0.107$\pm$0.030 0.080$\pm$0.045 0.036$\pm$0.032 0.026$\pm$0.020 0.253$\pm$0.166 0.076$\pm$0.088 0.000$\pm$0.000 0.000$\pm$0.000
GPTSelect 0.913$\pm$0.041 0.913$\pm$0.041 0.913$\pm$0.041 0.913$\pm$0.041 0.065$\pm$0.053 0.146$\pm$0.179 0.077$\pm$0.088 0.020$\pm$0.028
ATK 0.970$\pm$0.024 0.970$\pm$0.024 0.970$\pm$0.024 0.970$\pm$0.024 0.903$\pm$0.028 0.903$\pm$0.028 0.903$\pm$0.028 0.903$\pm$0.028
Table 3: Simulator Policy Success rate using different keypoint selection methods over 3 random seeds. $ATK$ consistently outperforms alternative keypoint selection methods using random sampling or ChatGPT selection.
@@ E.2 Imitation learning performance
Table 5: Real World Imitation Learning Success rate using different keypoint selection methods. $ATK$ consistently outperforms alternative keypoint selection methods using random sampling or ChatGPT selection.
@@
Learning rate $1\times 10^{-5}$ $1\times 10^{-5}$$1\times 10^{-4}$
Batch size $4096$ 20482048204820484096409640964096
Sparsity Weight $\lambda$ $0.002$ 0.0010.0010.0010.0010.0050.0050.0050.005

### v2

@@ A.1 Mask network
To ensure the selection of the minimal set of $K$ keypoints, we enforce a sparsity penalty on the mask network. We denote the masking distribution over binary masks $m\in\{0,1\}^{N}$ for candidate keypoints $K$ as $\mathbb{M}_{\phi}(m\mid K),$ and the action distribution conditioned on the masked keypoints as $\pi_{\theta}\bigl(a\mid\mathbb{M}_{\phi}).$ Because sampling $m\sim\mathbb{M}_{\phi}(m\mid K)$ is discrete and non-differentiable, we employ the Gumbel–softmax relaxation [22] to enable gradient-based optimization. Concretely, a $N$-dimensional learnable parameters are used as logits for N independent Bernoulli probabilities, with each representing the likelihood of retaining the corresponding keypoint from the input candidate set. We then add Gumbel noise to each logit, divide by a temperature $\tau$, and apply softmax to form a continuous “soft" mask $y^{Soft}_{M}$. During the forward pass, we take $\arg\max_{i}\,y^{\mathrm{Soft}}_{i},$ to get a hard one-hot selection, but in the backward pass we use the smooth “soft" mask for gradient calculation. The policy $\pi_{\theta}$ can be instantiated by any policy class model—e.g., a multi-layer perceptron, a Gaussian policy [46] or a score-based diffusion model [47]. The final objective jointly optimizes the policy $\pi_{\theta}(a\mid\mathbb{M}_{\phi})$ and the masking distribution $\mathbb{M}_{\phi}(m\mid K)$, yielding a minimal, task-relevant keypoint representation.
@@ A.2 Viewpoint robustness
A.3 Category-Level Generalization
To evaluate whether this method can generalize to different objects within the same category, we tested it on two tasks as shown in Figure 9. We assume that simulated objects may differ in color, size, or shape from their real-world counterparts; as long as they share the same category, a web-scale pretrained correspondence model can still match them reliably across various visual domain shifts. To measure the robustness of the correspondence matching algorithm, we consider two metrics: (1) Confidence Score, the mean cosine similarity of matched feature vectors, and (2) Mean Distance Error, the average pixel offset between manual annotations and the correspondences predicted by $h_{C}$. As shown in the Table 1 , the correspondence matching method achieves a high confidence score (0.76–0.82) and a low distance error (6 pixels, < 5% of the object size), demonstrating accurate sim-to-real keypoint transfer.
Figure 9: Category Level Generalization
Method Sushi Glass Clock Button Clock Turning
Confidence Score 0.78 0.76 0.80 0.82
Mean Distance Error (pixels) 3.24 5.21 2.79 6.75
Table 1: Correspondence Matching Performance Metrics
A.4 High-precision task
Mask model $\mathbb{M}_{\phi}$, keypoint policy $\pi_{\theta}$
// Initialize keypoints on canonical frame
Choose canonical image $I_{\text{can}}$ capturing full task context
Sample $C$ candidate keypoints $K_{\text{can}}\!=\!\{k_{j}\}_{j=1}^{C}$ on $I_{\text{can}}$
// Propagate & label entire dataset
for $i\leftarrow 0$ to M1M-1 do $\triangleright$ for all trajectories
  for $t\leftarrow 0$ to $T_{i}-1$ do
   if $t=0$ then
     $K_{i,0}\leftarrow h_{\mathcal{C}}\!\bigl(K_{\text{can}},\,o_{i,0},\,I_{\text{can}}\bigr)$ $\triangleright$ initial alignment to canonical frame
   else
     $K_{i,t}\leftarrow h_{\mathcal{C}}\!\bigl(K_{i,t-1},\,o_{i,t}\bigr)$ $\triangleright$ track from previous frame
   end if
  end for
end for
Construct $\mathcal{D}_{k}=\{(o_{i,t},K_{i,t},a_{i})\}$
//Training
while not converged do
  Sample mini-batch $\mathcal{B}\subset\mathcal{D}_{k}$
  Draw mask $m\sim\mathbb{M}_{\phi}(K)$ $\triangleright$ Gumbel–Softmax
  $\widetilde{K}\leftarrow m\odot K$ $\triangleright$ retain selected points
  Update $\phi,\theta$ with Adam to minimise $\mathcal{L}$
end while
return $(\mathbb{M}_{\phi},\pi_{\theta})$
function Evaluation
  // Transfer keypoints to test scene
  $K_{0}\leftarrow h_{\mathcal{C}}\!\bigl(K_{\text{can}},\,I_{\text{can}},\,O_{init}\bigr)$ $\triangleright$ highest-score matches
  Track $K_{t}=h_{\mathcal{C}}(K_{t-1},o_{t})$ for $t\geq 1$
  return $\pi_{\theta}(K_{t})$ at each step
end function
@@ Appendix D Baseline implementation details
ResNet18 Encoder: We use a modified ResNet18 architecture as the encoder for depth inputs. The depth data is first resized to $\times224$ and duplicated across three channels to match the expected input format of the ResNet18 backbone. Then, we take the ResNet’s 512-dimensional feature embedding as a feature map and concatenate it with a 7-dimensional robot state (six joint values and a binary gripper state). The resulting vector is then passed to the diffusion policy [47] for action prediction.
[…]
@@
Baseline Performance Notes: All baselines (RGB, depth, point-cloud) first used Appendix E.1 encoders with an MLP policy, but performance was weak, so we switched to the official diffusion-policy. With the same 50–60 demos as ATK, each baseline achieved >0% success on at least one task. However, visual domain shifts and sensor noise especially for depth and pointcloud caused failures, whereas ATK remained robust.
@@ E.1 Sim-to-real performance
Table 3: Real-world Policy Success Rates. Varying conditions including RP (random pose), RB (background), RO (distractor object), Light. ATKATK consistently outperforms baseline methods using alternative modalities in sim-to-real transfer.
@@
[…]
Table 4: Simulator Policy Success rate using different keypoint selection methods over 3 random seeds. ATKATK consistently outperforms alternative keypoint selection methods using random sampling or ChatGPT selection.
@@ E.2 Imitation learning performance
Table 6: Real World Imitation Learning Success rate using different keypoint selection methods. ATKATK consistently outperforms alternative keypoint selection methods using random sampling or ChatGPT selection.
@@
Learning rate $\times 10^{-5} 1×1$$\times 10^{-5}$$1\times 10^{-4}$
Batch size 40964096 2048204840964096
Sparsity Weight $\lambda$ $0.002$ 0.0010.0010.0050.005
@@ E.3 GPT selection details
E.4 Additional KeyPoint Selection Baselines
Figure 15: First row: GPT results when prompted with the initial set of candidate keypoints. Second row: RoboTAP selection results. Different colors denote distinct clusters of keypoints remaining after selection.
To make fairer comparison, we provide GPT with the same set of candidate keypoints as used for the initial set of keypoints on the canonical image from ATK. As illustrated in the figure 15 , the GPT-Select approach occasionally suffers from spatial mismatches (incorrectly locating the relevant point in the image) or language mismatches (incorrectly interpreting the textual prompt), which leads to erroneous keypoint selections. We also evaluated Robotap[53] keypoint’s selector on our tasks. As shown in the above figure, its design for ego-centric images prevents it from filtering out irrelevant points effectively in our setup.